175mL of water are added to 25 mL of a solution of HCl that pH =2.1. 

Determine the new value of the pH.


1
Expert's answer
2021-06-01T02:45:24-0400

Molar Mass of H2O = 18.01528

= 175/18.01528

= 9.714 moles

Molar Mass of HCl = 36.458

= 25/36.458

= 0.686 moles

= 9.714/0.686

= 14.16

Log 14.16

= 1.151

= 1.151 + 2.1

= 3.251